Metronidazole is used in the treatment of anaerobic protozoal infections. Although often administered orally, it can also be delivered by means of an IV infusion. The loading dose is 15 mg per kilogram of body mass and is infused over one hour. The maintenance dose is 7.5 mg per kilogram of body mass, is also infused over an hour, and is administered six hours after the loading dose. If the concentration of the metronidazole solution is 5 mg/ml, how many milliliters are required for a loading dose and three maintenance doses for a patient who weighs 220 pounds?
Answer #1

. 220pounds weight equal to 100kg

Tatal metronidazole given patient = (15mg+7.5×3mg)×100

15++22.5=37.5mg /kg

For 100 kg= 37.5×100=3750mg

Total solution required= 3750/5=750ml solution of 5mg/ml solution of metronidazole

Answer 750ml